OpenStudy (anonymous):

Myra's bulletin board has a width of 60 inches. When she originally decorated it she needed 192 inches of border material. How many inches high is the bulletin board?

OpenStudy (anonymous):

192 is the perimeter. Perimeter is calculated by: 2w + 2l P = 2(60) + 2l 192 = 120 + 2l 72 = 2l l = 36 inches

OpenStudy (anonymous):

The largest swimming pool in the world is in Casablanca. The rectangular pool has a perimeter of 1110 meters, and its length is 30 meters longer than 6 times its width. How many meters wide is it?